OpenStudy (anonymous):

slope 5/6, y-intercept (0,1/3) Express answer in standard form

OpenStudy (anonymous):

the standard form is: y=mx+b you got slope m=5/6---->y=5/6x+b y-int (0,1/3) plug this point in the equation: x=0, y=1/3---->1/3=5/6(0)+b 1/3=b therefore standard form of eqution is: y=5/6x+1/3

OpenStudy (anonymous):

Thank you so so much, I am struggling

OpenStudy (anonymous):

U r wlcm:)